OpenStudy (anonymous):

1+ cos(12x)=___ fill in the blank please!!!

OpenStudy (solomonzelman):

you want to re-write the expression (most likely).

OpenStudy (solomonzelman):

\(\large\color{slate}{ \cos(12x)~~\Rightarrow~~\cos^2(6x) -\sin^2(6x) }\) (double angle) \(\large\color{slate}{ 1-\cos(12x)~~\Rightarrow~~1+\cos^2(6x) -\sin^2(6x) }\) \(\large\color{slate}{ \Rightarrow~~1-\sin^2(6x) +\cos^2(6x) }\) \(\large\color{slate}{ \Rightarrow~~\cos^2(6x) +\cos^2(6x) }\) \(\large\color{slate}{ \Rightarrow~~2\cos^2(6x). }\)

OpenStudy (solomonzelman):

this is what I would do... but I don't know what they are exactly asking you here.

OpenStudy (anonymous):

That's exactly what they were looking for! Thank you!

OpenStudy (solomonzelman):

oh,. really? yw then...

OpenStudy (solomonzelman):

I just though you needed to get rid of the 1.... and yup I was right:)